OpenStudy (anonymous):

help please :] (6+3√2)(5-√2) ?? answers: 24+21√2 24+9√2 30+6√2 30+11√2

OpenStudy (anonymous):

Try using FOIL, what happens?

OpenStudy (anonymous):

i dont know how to use it on this.

OpenStudy (anonymous):

It's similar.

OpenStudy (anonymous):

Start with First. What do you get?

OpenStudy (anonymous):

try 1/6 [ (6+3√2)(5-√2)] multiply by reciprocal of 6

OpenStudy (anonymous):

simplifies to 3√2( 5-√2/6)

OpenStudy (anonymous):

can you finish the rest?

OpenStudy (anonymous):

keep in mind that √x * √x = x

OpenStudy (anonymous):

Then do outer: 6* -√2 and inner: 3√2 *5. add the like terms together.

OpenStudy (anonymous):

doh i was going to let skate try :P

OpenStudy (anonymous):

I'm not finishing it for him :P but I do remember radicals giving me trouble

OpenStudy (anonymous):

so ur adding 6 and-√2 together? im sorry im new at this and its freaking hard haha

OpenStudy (anonymous):

no

OpenStudy (anonymous):

dam

OpenStudy (anonymous):

so i will go over this again

OpenStudy (anonymous):

you have this (6+3√2) (5-√2) <-------- this is given

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

simplify 1/6 [ 6+3√2)(5-√2) ] multiply the whole equation by reciprocal of 6 ^ L____________ this is the reciprocal of 6 (it is 1/6) the number multiplied by its reciprocal will always = 1 , x * 1/x = 1 and -x * -1/x =1

OpenStudy (anonymous):

...I disagree, mth3v4. You can't divide the entire problem by 6. It's in the parentheses. If there were multiples of 6 in each term, then you could.

OpenStudy (anonymous):

mmmm okies

OpenStudy (anonymous):

you got a point

OpenStudy (anonymous):

careless mistake xD

OpenStudy (anonymous):

No worries, we've all done that :)

OpenStudy (anonymous):

OK, Skate, you have experience doing FOIL, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but can you factor the 6 out?

OpenStudy (anonymous):

Ok, so which terms are first, outer, inner, and last?

OpenStudy (anonymous):

Wait no u still get a 1 :P

OpenStudy (anonymous):

No, you can't.

OpenStudy (anonymous):

nm i didnt say nething

OpenStudy (anonymous):

You aren't factoring this problem, it's already factored.

OpenStudy (anonymous):

So, skate, what are the first terms?

OpenStudy (anonymous):

i was just saying factor out a 6 of the first parenthesis ( but you end w/ a 1+3√2< not what i want)

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

that's the first term in the first parentheses. What's the one in the second set?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

ok, so multiply those two together. What do you get?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

nice job! OK, so that's first. Now, what are the two outer terms?

OpenStudy (anonymous):

√2 and -√2 right?

OpenStudy (anonymous):

no, those are the Last terms, but we can do that now if you'd like

OpenStudy (anonymous):

Remember to keep the 3 in there though.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ohh alright

OpenStudy (anonymous):

well you can have a fun way to do FOIL :D first inner outer last but _____________ | |-----|----------| | ( 1x + 1y ) ( 2x + 2y ) \ L_______/ / \________________/ an easier(fun way to do FOIL (i prefer) make a happy face :)

OpenStudy (anonymous):

I like the happy face way!!

OpenStudy (anonymous):

teacher taught me not my credit

OpenStudy (anonymous):

So what do you get when you multiply the outer terms?

OpenStudy (anonymous):

its easier to keep track of what im doing

OpenStudy (anonymous):

you multiply 3 and -2 or 3 and √2? and thanks for the happy face foil method haha :]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

fun to be happy

OpenStudy (anonymous):

yup especially in math lol xP

OpenStudy (anonymous):

math is cool (when your good at it) xD

OpenStudy (anonymous):

for the last terms, you're right, it becomes 3 times -2

OpenStudy (anonymous):

ok so -6

OpenStudy (anonymous):

ok, so add -6 to 30. what do you get?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yep! So that narrows it down to two.

OpenStudy (anonymous):

Now, what are the outer terms and the inner terms?

OpenStudy (anonymous):

wait lemme check again

OpenStudy (anonymous):

i think what we are dealing with here is conjugate multiplication right?

OpenStudy (anonymous):

umm? im really not sure :p

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I don't actually know, sorry.

OpenStudy (anonymous):

ya its conjugate factored difference of squares

OpenStudy (anonymous):

Ok, skate, what are the outer terms?

OpenStudy (anonymous):

24? or is it 30?

OpenStudy (anonymous):

you got difference of squares with 3 √2, √2

OpenStudy (anonymous):

you already added the first and last terms together, and got 24

OpenStudy (anonymous):

its the squareroot

OpenStudy (anonymous):

we're trying to get the second part of the answer, which you need the outer and inner terms for

OpenStudy (anonymous):

√x *√ x = x

OpenStudy (anonymous):

um. that's what we did for the last terms. but that's not in the outer and inner terms.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Skate, using the happy face method that mth3v4 showed you, what are the outer terms? and then multiply them together.

OpenStudy (anonymous):

ohh so 2 and -2?

OpenStudy (anonymous):

wait never mind thats wrong

OpenStudy (anonymous):

so no more radical its just 30-3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Ok, so the outer terms are 6 and -√2. Multiply them together

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

its a difference of squares problem (1-√x^2)(1+√x^2) , (1 + x^2) (1- x^2) right? or is it another careless mistake :P (i actually make a lot of them, really )

OpenStudy (anonymous):

hah i knew it

OpenStudy (anonymous):

@Skate: not exactly. Replace √2 with x. What's 6 * x?

OpenStudy (anonymous):

@mth3v4: no, it's not a difference of squares problem

OpenStudy (anonymous):

okies i got 24

OpenStudy (anonymous):

6x?

OpenStudy (anonymous):

yep! so that makes it 6√2, right?

OpenStudy (anonymous):

so now the inner terms. can you find them?

OpenStudy (anonymous):

but i still think this is very much in the form a diff squares prob reasons because the expression was in the form of this (n + √x) (n-√x) n= whatever number

OpenStudy (anonymous):

diff of squares. (x+1)(x-1)= x^2

OpenStudy (anonymous):

radical √x *√x = x

OpenStudy (anonymous):

It is in the form of a difference of square problem in that one is positive and one is negative, but it doesn't fall under the definition of a difference of squares. Because there are no squares.

OpenStudy (anonymous):

OK, skate, inner terms are 3√2 * 5. multiply together, k?

OpenStudy (anonymous):

ok